# A Systematic Investigation of Knowledge Retrieval and Selection for Retrieval Augmented Generation

### v1

@@ Abstract
Retrieval-augmented generation (RAG) has emerged as a powerful method for enhancing natural language generation by integrating external knowledge into a model’s output. While prior work has demonstrated the importance of improving knowledge retrieval for boosting generation quality, the role of knowledge selection remains less clear. In this paper, we perform a comprehensive analysis of how knowledge retrieval and selection influence downstream generation performance in RAG systems. By simulating different retrieval and selection conditions through a controlled mixture of gold and distractor knowledge, we assess the impact of these factors on generation outcomes. Our findings indicate that the downstream generator model’s capability, as well as the complexity of the task and dataset, significantly influence the impact of knowledge retrieval and selection on the overall RAG system performance. In typical scenarios, improving the knowledge recall score is key to enhancing generation outcomes, with the knowledge selector providing a limited additional benefit when a strong generator model is used on clear, well-defined tasks. For weaker generator models or more ambiguous tasks and datasets, the knowledge F1 score becomes a critical factor, and the knowledge selector plays a more prominent role in improving overall performance.
A Systematic Investigation of Knowledge Retrieval and Selection for Retrieval Augmented Generation
Xiangci Li1, 2    Jessica Ouyang1 1 University of Texas at Dallas, 2 Amazon Web Services lixiangci8@gmail.com, jessica.ouyang@utdallas.edu
**footnotetext:  Work performed outside of AWS.
@@ 1 Introduction
The quality of the retrieved knowledge directly influences the quality of RAG outputs. Previous studies have consistently shown that improving knowledge retrieval leads to a direct enhancement in generation performance Dinan et al. (2019); Li et al. (2022, 2024); Cuconasu et al. (2024). Similarly, effective knowledge selection has been observed to improve generation quality by filtering out retrieved information that is less relevant to the generation target Kim et al. (2020); Thulke et al. (2021); Li et al. (2022); Sun et al. (2023); Zhang et al. (2023); Zheng et al. (2024). However, we observe that knowledge selection is rarely used for tasks other than dialogue generation, and there are notably fewer LLM-based RAG works that use knowledge selection, compared to fine-tuned RAG models (see Section 2). We hypothesize that knowledge selectors may not always improve downstream generation performance, and that there may be a selection bias where only positive results involving knowledge selector modules are published, while experiments where knowledge selectors are not helpful simply do not report those results.
In this work, we perform a systematic analysis of how knowledge retrieval and selection affect downstream RAG performance. By blending gold knowledge with distractor knowledge in varying ratios, we simulate different retrieval and selection outcomes and examine their impact on the overall performance of RAG. We find that the generator model’s capability, as well as the complexity of the task and dataset, significantly influence RAG system performance. In typical scenarios, improving knowledge recall via the knowledge retriever is key to enhancing generation outcomes, with the knowledge selector providing limited additional benefit when a strong generator model is used on clear, well-defined tasks. For weaker generator models or more ambiguous tasks and datasets, knowledge F1 becomes a critical factor, and the knowledge selector plays a more prominent role in improving overall performance.
Figure 1: Our simulation experiment pipeline.
@@ 2 Related Work
Knowledge-grounded dialogue generation
Moghe et al. (2018); Dinan et al. (2019); Li et al. (2024) is one of the major applications of RAG, where the target response is conditioned on retrieved knowledge, and a knowledge selection step is often added to further refine the retrieved knowledge Thulke et al. (2021); Sun et al. (2023); Zhang et al. (2023). For example, Kim et al. (2020) train a knowledge selector by leveraging response information, Li et al. (2022) select knowledge from document semantic graphs, and Zhang et al. (2023) propose multi-task learning for knowledge selection and response generation. However, despite these works demonstrating the advantages of their overall approaches through ablation studies, it is unclear whether their performance improvement via knowledge selection specifically is transferable to other datasets, domains, or tasks. Moreover, despite the popularity of LLMs for RAG, there are dramatically fewer LLM-based RAG works that include the knowledge selection step, and it is likewise omitted in other RAG tasks Gao et al. (2023); Fan et al. (2024).
To the best of our knowledge, there is no prior work explaining this gap in knowledge selection. The closest work to ours is Cuconasu et al. (2024) which studies the impact of knowledge retrieval on RAG. We hypothesize that there is a selection bias effect where knowledge selection may have been experimented with, but ultimately not reported, in tasks or settings where it is not helpful. In this paper, we study the general effect of knowledge retrieval and selection on LLM-based RAG via simulations, rather than being restricted to specific retriever or selector implementations.
@@ 3.2 Knowledge Simulation
We aim to perform a systematic analysis of the effect of knowledge retrieval and selection outcomes on the downstream RAG performance by simulation. As Figure 1 shows, for each query, given a fixed pool of available knowledge with gold knowledge annotations, we sample gold knowledge and distractor knowledge at varying rates $p_{gold}$ and $p_{noise}$ to precisely simulate a wide range of quality for the retrieved and selected knowledge $K^{\prime}$, from noise only to gold knowledge only. For example, if $p_{gold}=0.5$, then each piece of gold knowledge has a 50% chance to be sampled. In this way, we simulate a wide distribution of knowledge retrievers and selectors of different performance levels, as measured by knowledge precision and recall, which is computed based on the annotated gold knowledge. We then test the performance of the generator model given the different quality of retrieved and selected knowledge $K^{\prime}$.
@@ 4.1 Datasets
is an open-domain dialogue dataset based on Wikipedia knowledge that has been widely used in prior knowledge selection works Kim et al. (2020); Thulke et al. (2021); Li et al. (2022); Sun et al. (2023). The wizard speaker (a human annotator playing the part of the dialogue system) has access to Wikipedia knowledge, while the apprentice speaker does not. The apprentice is given a starting conversation topic, but otherwise speaks freely. The last two turns of dialogue are used as the query to retrieve Wikipedia passages ($K$), and the wizard selects a single knowledge sentence ($K^{\prime}$) to generate their response.
Despite its popularity, this dialogue generation task is challenging to evaluate. First, because the wizard is limited to selecting only a single knowledge sentence, it is possible for the unselected “distractor" knowledge to also be relevant to the wizard’s response. Second, as is the case in many natural language generation tasks, the gold responses in WoW are not the only plausible responses, making it harder to quantify the correctness of a generated response. Nonetheless, WoW is a well-annotated dialogue dataset that is close to a real-life scenario where the gold knowledge and responses are noisy.
@@
is a question-answering dataset derived from Wikipedia knowledge that we include in our experiments to mitigate the ambiguity of WoW. The multi-hop questions and answers are first derived from gold knowledge graphs, and then distractor knowledge is injected to create noise. As a result, unlike WoW, the answers in HotpotQA are strongly dependent on the gold knowledge, and the distractor knowledge is unlikely to be relevant. The short and unambiguous gold answers make it simpler to evaluate via F1 scores.
@@ 4.2 Generators
To align with the latest trends in RAG research, as well as to make the analysis simple and generally replicable, we adopt LLMs as our generators. We use three API-based LLMs, OpenAI GPT-4o-mini, LLaMA 3.1 8B, and Mistral 7B-Instruct, as our generator models; the varying sizes of the LLMs allow us to investigate the impact of generator complexity on knowledge usage.
@@
No knowledge 0 0 0 0.110 0.200
Full knowledge 0.015 1 0.031 0.140 0.251
Gold knowledge 1 1 1 0.167 0.276
LLM selector 0.054 0.720 0.101 0.139 0.247
@@
No knowledge 0 0 0 0.111 0.216
Full knowledge 0.015 1 0.031 0.138 0.248
Gold knowledge 1 1 1 0.164 0.278
LLM selector 0.053 0.317 0.009 0.126 0.231
@@
No knowledge 0 0 0 0.113 0.203
Full knowledge 0. 1 0. 0.131 0.233
Gold knowledge 1 1 1 0.172 0.268
LLM selector 0.036 0.522 0.067 0.123 0.221
Table 1: WoW response generation performance benchmarked by different LLM generatorss. We measure knowledge precision (KP), recall (KR), and F1 (KF1); response ROUGE-L F1 (R-L); and response F1.
@@
No knowledge 0 0 0 0.330 0.437
Full knowledge 0.059 1 0.111 0.668 0.780
Gold knowledge 1 1 1 0.710 0.828
LLM selector 0.351 0.685 0.464 0.622 0.724
@@
No knowledge 0 0 0 0.200 0.298
Full knowledge 0.059 1 0.111 0.372 0.545
Gold knowledge 1 1 1 0.414 0.671
LLM selector 0.304 0.405 0.347 0.294 0.469
@@
No knowledge 0 0 0 0.208 0.260
Full knowledge 0.059 1 0.046 0.668 0.151
Gold knowledge 1 1 1 0.502 0.627
LLM selector 0.06 0.244 0.103 0.186 0.245
Table 2: HotpotQA answer generation performance benchmarked by different LLM generators. We measure knowledge precision (KP), recall (KR), and F1 (KF1); answer exact match (EM); and answer F1.
@@ 4.3 Knowledge Sampling
Since knowledge precision and recall are the most common metrics for knowledge retrieval and selection performance, we use these metrics as the basis for our analysis. We find that $p_{gold}$ linearly correlates with knowledge recall scores, while $p_{noise}$ exponentially correlates with knowledge precision. Thus, to simulate retrievers/selectors with specific knowledge precision and recall scores, we use grid search in the linear space of $p_{gold}$ and both the linear and exponential spaces of $p_{noise}$ to ensure most grids in the knowledge precision-recall space are covered by our experiments.
As Figure 1 shows, both WoW and HotpotQA provide a retrieved knowledge set $K$ for each query $q$ such that the gold knowledge is a subset of $K$. In both datasets, $K$ is retrieved using bigram TF*IDF Chen et al. (2017) and contains text passages, where each passage consists of an article title and a few knowledge sentences. We perform knowledge sampling at the sentence level; the sampled knowledge simulates the end result of applying both the knowledge retrieval and selection steps.
In our experiments, we refer to using the entire provided knowledge set $K$ as the input knowledge to the generator (i.e. no knowledge selection) as the “full knowledge” setting. We also include the “no knowledge" setting, where the generators receive no external knowledge at all, as a weak baseline, and “gold knowledge," corresponding to perfect knowledge selection, as an upper bound.
Figure 2: Scatter plot of response/answer F1, plotted against knowledge precision (x-axis) and recall (y-axis), by GPT-4o-mini (top), LLaMA 3.1 8B (middle), and Mistral 7B-Instruct (bottom). The left column shows results on WoW; the right shows HotpotQA. The dots highlighted in orange indicate settings outperforming the “full knowledge” setting, while those highlighted in cyan indicate settings underperforming the “no knowledge" setting.
4.4 Implementation Details
We do not perform any fine-tuning or hyperparameter tuning. We set the LLMs’ temperature to 0 and ensure that our zero-shot generation prompts (Appendix A.1) are shorter than their maximum input lengths. We use the API services of OpenAI gpt-4o-mini-2024-07-18***https://platform.openai.com/docs/models/gpt-4o-mini, Together AI***https://docs.together.ai/docs/chat-models meta-llama/Meta-Llama-3.1-8B-Instruct-Turbo, and mistralai/Mistral-7B-Instruct-v0.1. We report results on the first 500 examples in the HotpotQA training set and the first 100 conversations (452 wizard utterances) from the “test seen” set of WoW.
@@ 5 Simulation Results
Overall we observe several consistent trends, except for Mistral-7B-Instruct, which is a relatively weak-performing LLM.
@@
As Tables 1 & 2 show, generators without any retrieved knowledge (“no knowledge" setting) perform poorly on WoW and HotpotQA, even though they may have pre-trained on the Wikipedia articles that WoW and HotpotQA are derived from. This finding indicates the LLMs are not over-fitted on the WoW and HotpotQA datasets, and applying RAG is beneficial.
Interestingly, our HotpotQA results in Figure 2 show that distractor knowledge significantly harms performance; the cyan dots show that generators receiving mostly distractor knowledge underperform the “no knowledge” setting, which uses only the LLM’s internal knowledge. This trend is not observed with WoW, supporting our observation that HotpotQA’s distractor knowledge is truly irrelevant, while the “distractor" knowledge in WoW may only be less relevant than the single gold sentence, but not completely irrelevant.
Figure 3: Scatter plot of HotpotQA answer F1 versus knowledge precision (top), knowledge recall (middle), and knowledge F1 (bottom). The left column shows GPT-4o-mini as the generator; the right column shows Mistral-7B-Instruct. Plots for LLaMa 3.1 8B and the WoW dataset are in Appendix A.
@@
Tables 1 & 2 show that the “full knowledge" setting, which corresponds to knowledge retrieval with perfect recall, but no knowledge selection, is a very strong baseline. The tables also include an LLM knowledge selector, implemented with the same LLM as the generator (prompts in Appendix A.1), to show the performance of a typical knowledge selector module. None of these LLM knowledge selectors outperform the “full knowledge” setting, except Mistral-7B-Instruct on HotpotQA.
This finding mostly contrasts with those of prior knowledge selection works Kim et al. (2020); Thulke et al. (2021); Li et al. (2022); Sun et al. (2023); Zhang et al. (2023); Zheng et al. (2024). For example, in Table 2, we find that for GPT-4o-mini, a very strong generator model, the “full knowledge" setting achieves 0.780 answer F1 on HotpotQA, only 0.048 lower than the “gold knowledge" F1; similar observations are seen for LLaMA 3.1 8B and for both models on WoW. Since the performance gap between the “full knowledge” and “gold knowledge” settings is the space for a knowledge selector to improve generation performance, we conclude that strong generators simply have less room for improvement via knowledge selection.
Knowledge precision & recall together are good indicators of generation performance.
Figure 2 shows that generation performance varies smoothly with knowledge precision and recall, indicating that knowledge precision and recall together are major determinants of generation performance. Due to the pipeline nature of the retriever, selector, and generator components of RAG, knowledge recall can only be improved by the retriever. The selector can only improve knowledge precision, but it is likely to also reduce recall. Therefore, we can visualize the effect of applying a knowledge selector: it moves the RAG system towards the lower-right corner of the plots in Figure 2; the better the knowledge selector, the more it will move right, and the less it will move down.
@@
We find that for strong generator models, the knowledge recall score is the best single knowledge metric for estimating generation performance; Figure 3 shows a very strong correlation between knowledge recall and answer F1 for GPT-4o-mini on HotpotQA, and Figure 8 in the Appendix shows the same for LLaMA 3.1 8B and both models on WoW. Moreover, knowledge recall is the most important factor in improving generation performance given a fixed generator, while the knowledge precision score is a secondary factor.
@@
Figure 4: Color contours of answer F1 versus the knowledge precision for GPT-4o-mini (top), LLaMA 3.1 8B (middle), and Mistral 7B-Instruct (bottom); the left column shows results on WoW, and the right shows HotpotQA. Each contour represents a different knowledge recall score; moving left to right visualizes improving the performance (precision) of the knowledge selector.
@@
All of our results show that the overall RAG performance depends on the generator model: the cross-model comparisons in Figures 2 & 4 show that for stronger generator models (as measured by rank in LLM leaderboards***https://huggingface.co/spaces/lmsys/chatbot-arena-leaderboard), performance across all knowledge settings are higher (Tables 1 & 2). While this result is unsurprising, we also find that the gap between the “full knowledge” and “gold knowledge” settings becomes narrower, suggesting that stronger generators are more robust to noisy input knowledge and rely less on the knowledge selector. In contrast, when the generator is weaker, any reasonable knowledge selector becomes beneficial, likely because a weak generator cannot handle noisy input knowledge and thus requires a selector to filter out distractor knowledge — knowledge F1 best correlates to the answer F1 for Mistral-7B-Instruct in Figure 3.
@@
In addition, Figure 4 shows that attempting to improve the knowledge selector on top of a weak retriever (i.e. increasing precision when recall is low) hurts generation performance on WoW, as well as for GPT-4o-mini on Hotpot QA. In this scenario, more total knowledge, regardless of noise, improves response generation. These counter-intuitive observations are likely due to the nature of the task and the annotation quality. The solution space in HotpotQA is small given a specific question, and HotpotQA has a cleaner separation between gold and distractor knowledge, whereas in WoW, there are more plausible responses given the same conversation history and knowledge, as well as “distractor" knowledge that may actually be relevant to the gold response.
@@
Interestingly, we observe that the boundary of where the knowledge selector improves generation performance (the border between the orange and white areas in Figure 2) is convex for all three LLMs on WoW. This phenomenon can also be seen in Figure 4, where some color contours intersect with the “full knowledge answer F1” baseline (dashed red line) multiple times. In other words, we observe that generation performance is non-monotonic as knowledge precision increases with fixed knowledge recall. The fact that this phenomenon is only observed on WoW may be due to the relatively noisy gold knowledge annotations in WoW; we can produce similar behavior by artificially injecting noise into HotpotQA’s gold knowledge annotations (Appendix A.3).

### v2

# How Does Knowledge Selection Help Retrieval Augmented Generation?

@@ Abstract
Retrieval-augmented generation (RAG) is a powerful method for enhancing natural language generation by integrating external knowledge into a model’s output. While prior work has demonstrated the importance of improving knowledge retrieval for boosting generation quality, the role of knowledge selection remains less clear. In this paper, we perform a comprehensive analysis of how knowledge selection influences downstream generation performance in RAG systems. By simulating different retrieval and selection conditions through a controlled mixture of gold and distractor knowledge, we assess the impact of these factors on generation outcomes. Our findings indicate that the downstream generator model’s capability, as well as the complexity of the task and dataset, significantly influence the impact of knowledge selection on the overall RAG system performance. In typical scenarios, improving the knowledge recall score is key to enhancing generation outcomes, with the knowledge selector providing a limited additional benefit when a strong generator model is used on clear, well-defined tasks. For weaker generator models or more ambiguous tasks and datasets, the knowledge F1 score becomes a critical factor, and the knowledge selector plays a more prominent role in improving overall performance.
How Does Knowledge Selection Help Retrieval Augmented Generation?
Xiangci Li1, 2    Jessica Ouyang2 1 AWS AI Labs, 2 University of Texas at Dallas lixiangci8@gmail.com, jessica.ouyang@utdallas.edu
**footnotetext:  Work performed outside of AWS AI Labs.
@@ 1 Introduction
The quality of the retrieved knowledge directly influences the quality of RAG outputs. Previous studies have consistently shown that improving knowledge retrieval leads to a direct enhancement in generation performance Dinan et al. (2019); Li et al. (2022, 2024); Cuconasu et al. (2024); Wang et al. (2024); Wu et al. (2024). Similarly, effective knowledge selection has been observed to improve generation quality by filtering out retrieved information that is less relevant to the generation target Kim et al. (2020); Thulke et al. (2021); Li et al. (2022); Sun et al. (2023); Zhang et al. (2023); Zheng et al. (2024); Wang et al. (2024). However, we observe that knowledge selection is rarely used for tasks other than dialogue generation, and there are notably fewer LLM-based RAG works that use knowledge selection, compared to fine-tuned RAG models (see Section 2). We hypothesize that knowledge selectors may not always improve downstream generation performance, and that there may be a selection bias where only positive results involving knowledge selector modules are published, while experiments where knowledge selectors are not helpful simply do not report those results.
In this work, we perform a systematic analysis of how knowledge selectors with various performances affect downstream RAG performance. By blending gold knowledge with distractor knowledge in varying ratios, we simulate different selection outcomes and examine their impact on the overall performance of RAG. We find that the generator model’s capability, as well as the complexity of the task and dataset, significantly influence RAG system performance. In typical scenarios, improving knowledge recall via the knowledge retriever is key to enhancing generation outcomes, with the knowledge selector providing limited additional benefit when a strong generator model is used on clear, well-defined tasks. For weaker generator models or more ambiguous tasks and datasets, knowledge F1 becomes a critical factor, and the knowledge selector plays a more prominent role in improving overall performance.
Figure 1: Our simulation experiment pipeline for each example.
@@ 2 Related Work
Knowledge selector for knowledge-grounded dialogue generation.
Dialogue generation Moghe et al. (2018); Dinan et al. (2019); Li et al. (2024) is one of the major applications of RAG, where the target response is conditioned on retrieved knowledge, and a knowledge selection step is often added to further refine the retrieved knowledge Thulke et al. (2021); Sun et al. (2023); Zhang et al. (2023). For example, Kim et al. (2020) train a knowledge selector by leveraging response information, Li et al. (2022) select knowledge from document semantic graphs, and Zhang et al. (2023) propose multi-task learning for knowledge selection and response generation. However, despite these works demonstrating the advantages of their overall approaches through ablation studies, it is unclear whether their performance improvement via knowledge selection specifically is transferable to other datasets, domains, or tasks. Moreover, despite the popularity of LLMs for RAG, there are dramatically fewer LLM-based RAG works that include the knowledge selection step, and it is likewise omitted in other RAG tasks Gao et al. (2023); Fan et al. (2024).
To the best of our knowledge, there is no prior work explaining this gap in knowledge selection. The closest work to ours are Cuconasu et al. (2024) and Wu et al. (2024), which study the impact of knowledge retrieval on RAG. We hypothesize that there is a selection bias effect where knowledge selection modules may have been experimented with, but ultimately not reported, in tasks or settings where it is not helpful. In this paper, we study the general effect of knowledge retrieval and selection on LLM-based RAG via simulations, rather than being restricted to specific retriever or selector implementations.
@@ 3.2 Knowledge Simulation
We aim to perform a systematic analysis of the effect of knowledge retrieval and selection outcomes on the downstream RAG performance by simulation. As Figure 1 shows, for each query $q$, given a fixed pool of available knowledge with gold relevance annotations, we sample gold knowledge and distractor knowledge at varying rates $p_{gold}$ and $p_{noise}$ to precisely simulate a wide range of quality for the retrieved and selected knowledge $K^{\prime}$, from noise only to gold knowledge only. For example, if $p_{gold}=0.5$, then each piece of gold knowledge has a 50% chance to be sampled. In this way, we simulate a wide distribution of knowledge retriever and selector performance, as measured by knowledge precision and recall computed based on the annotated gold knowledge. We then test the performance of the generator model given the different quality of simulated retrieved and selected knowledge $K^{\prime}$.
@@ 4.1 Datasets
is an open-domain dialogue dataset based on Wikipedia knowledge that has been widely used in prior knowledge selection works Kim et al. (2020); Thulke et al. (2021); Li et al. (2022); Sun et al. (2023). The wizard speaker (a human annotator playing the part of the dialogue system) has access to Wikipedia knowledge, while the apprentice speaker does not. The apprentice is given a starting conversation topic, but otherwise speaks freely. The last two turns of dialogue are used as the query $q$ to retrieve Wikipedia passages ($K$), and the wizard selects a single knowledge sentence ($K^{\prime}$) to generate their response.
Despite its popularity, this dialogue generation task is challenging to evaluate. First, because the wizard is limited to selecting only a single knowledge sentence, it is still possible for the unselected “distractor" knowledge to be relevant to the wizard’s response. Second, as is the case in many natural language generation tasks, the gold wizard responses in WoW are not the only plausible responses, making it harder to quantify the correctness of a generated response. Nonetheless, WoW is a well-annotated dialogue dataset that is close to a real-life scenario where the gold knowledge and responses are noisy.
@@
is a question-answering dataset derived from Wikipedia knowledge that we include in our experiments to mitigate the ambiguity of WoW. The multi-hop questions and answers are first directly derived from gold knowledge graphs, and then distractor knowledge is injected to create noise. As a result, unlike WoW, the answers in HotpotQA are strongly dependent on the gold knowledge, and the distractor knowledge is unlikely to be relevant. The short and unambiguous gold answers make it simpler to evaluate via F1 scores.
@@ 4.2 Generators
To align with the latest trends in RAG research, as well as to make the analysis simple and generally replicable, we adopt LLMs as our generators. We use three API-based LLMs, OpenAI GPT-4o-mini, LLaMA 3.1 8B, and Mistral 7B-Instruct, as our generator models; the varying performances of the LLMs allow us to investigate the impact of generator complexity on knowledge usage.
@@
No knowledge 0 0 0 0.110 0.200 ($\pm$ .005)
Full knowledge 0.015 1 0.031 0.140 0.251 ($\pm$ .006)
Gold knowledge 1 1 1 0.167 0.276 ($\pm$ .007)
@@
No knowledge 0 0 0 0.111 0.216 ($\pm$ 0.005)
Full knowledge 0.015 1 0.031 0.138 0.248 ($\pm$ .005)
Gold knowledge 1 1 1 0.164 0.278 ($\pm$ .008)
@@
No knowledge 0 0 0 0.113 0.203 ($\pm$ .005)
Full knowledge 0 1 0 0.131 0.233 ($\pm$ .005)
Gold knowledge 1 1 1 0.172 0.268 ($\pm$ .007)
Table 1: WoW response generation performance benchmarked by different LLM generators. We measure knowledge precision (KP), recall (KR), and F1 (KF1); response ROUGE-L F1 (R-L); and response F1 (and its standard error mean).
@@
No knowledge 0 0 0 0.330 0.437 ($\pm$ .020)
Full knowledge 0.059 1 0.111 0.668 0.780 ($\pm$ .016)
Gold knowledge 1 1 1 0.710 0.828 ($\pm$ .014)
@@
No knowledge 0 0 0 0.200 0.298 ($\pm$ .019)
Full knowledge 0.059 1 0.111 0.372 0.545 ($\pm$ .019)
Gold knowledge 1 1 1 0.414 0.671 ($\pm$ .016)
@@
No knowledge 0 0 0 0.208 0.260 ($\pm$ .019)
Full knowledge 0.059 1 0.046 0.668 0.151 ($\pm$ .011)
Gold knowledge 1 1 1 0.502 0.627 ($\pm$ .019)
Table 2: HotpotQA answer generation performance benchmarked by different LLM generators. We measure knowledge precision (KP), recall (KR), and F1 (KF1); answer exact match (EM); and answer F1 (and its standard error mean).
@@ 4.3 Knowledge Sampling
Since knowledge precision and recall are the most common metrics for knowledge retrieval and selection performance, we use these metrics as the basis for our analysis. We find that the sampling rate $p_{gold}$ linearly correlates with knowledge recall scores, while $p_{noise}$ exponentially correlates with knowledge precision. Thus, to simulate retrieving and selecting a knowledge set $K^{\prime}$ with a specific knowledge precision and recall score, we use grid search in the linear space of $p_{gold}$ and both the linear and exponential spaces of $p_{noise}$ to ensure most grids in the knowledge precision-recall space are covered by our experiments.
As Figure 1 shows, both WoW and HotpotQA are knowledge selector training datasets that provide a retrieved knowledge set $K$ for each query $q$ such that the gold knowledge is a subset of $K$. In both datasets, $K$ contains text passages consisting of an article title and a few knowledge sentences. We perform knowledge sampling at the sentence level to simulate the end result of applying both the knowledge retrieval and selection steps.
In our experiments, we refer to using the entire retrieved knowledge set $K$ (provided by the dataset) as the input knowledge to the generator (i.e. no knowledge selection) as the “full knowledge” setting, where both gold and distractor knowledge are present. We also include the “no knowledge" setting, where the generators receive no external knowledge at all, as a weak baseline, and “gold knowledge," corresponding to perfect knowledge selection, as an upper bound.
Figure 2: Scatter plot of response/answer F1, plotted against knowledge precision (x-axis) and recall (y-axis), by GPT-4o-mini (top), LLaMA 3.1 8B (middle), and Mistral 7B-Instruct (bottom). The left column shows results on WoW; the right shows HotpotQA. The dots highlighted in orange indicate settings outperforming the “full knowledge” setting, while those highlighted in cyan indicate settings underperforming the “no knowledge" setting. Each dot corresponds to a full experiment on the entire sampled dataset.
@@ 5 Simulation Results
Overall we observe several consistent trends, except for Mistral-7B-Instruct, which is a relatively weak-performing LLM. Note that each point plotted in Figures 2-4 corresponds to a full experiment on the entire sampled dataset. The standard error mean values in Tables 1 & 2 indicate that the answer F1 scores are robust across all experiments and the overall observed trends are significant.
@@
As Tables 1 & 2 show, generators without any retrieved knowledge (“no knowledge" setting) perform poorly on WoW and HotpotQA, even though they may have been pre-trained on the Wikipedia articles that WoW and HotpotQA are derived from. This finding indicates the LLMs are not over-fitted on the WoW and HotpotQA datasets, and applying RAG is beneficial.
Interestingly, our HotpotQA results in Figure 2 show that distractor knowledge significantly harms performance; the cyan dots show that generators receiving mostly distractor knowledge underperform the “no knowledge” setting, which uses only the LLM’s internal knowledge. This trend is not observed with WoW, supporting our observation that HotpotQA’s distractor knowledge is truly irrelevant, while the “distractor" knowledge in WoW may only be less relevant than the single gold sentence, but not completely irrelevant. This difference between the two datasets is further explored in Appendix A.4.
Figure 3: Scatter plot of HotpotQA answer F1 versus knowledge precision (top), knowledge recall (middle), and knowledge F1 (bottom). The left column shows GPT-4o-mini as the generator; the right column shows Mistral-7B-Instruct. Plots for LLaMa 3.1 8B and the WoW dataset are in Appendix A. Each point corresponds to a full experiment on the entire sampled dataset.
@@
Tables 1 & 2 show that the “full knowledge" setting, which corresponds to knowledge retrieval with perfect recall, but no knowledge selection, is a very strong baseline.
This finding mostly contrasts with those of prior knowledge selection works Kim et al. (2020); Thulke et al. (2021); Li et al. (2022); Sun et al. (2023); Zhang et al. (2023); Zheng et al. (2024); Wang et al. (2024). For example, in Table 2, we find that for GPT-4o-mini, a very strong generator model, the “full knowledge" setting achieves 0.780 answer F1 on HotpotQA, only 0.048 lower than the “gold knowledge" setting; similar observations are seen for LLaMA 3.1 8B on HotpotQA and for both models on WoW. Since the performance gap between the “full knowledge” and “gold knowledge” settings is the space for a knowledge selector to improve generation performance, we conclude that strong generators simply have less room for improvement via knowledge selection.
Knowledge precision & recall together are good predictors of generation performance.
Figure 2 shows that generation performance varies smoothly with knowledge precision and recall, indicating that knowledge precision and recall together are major determinants of generation performance. Due to the pipeline nature of the retriever, selector, and generator components of RAG, knowledge recall can only be improved by the retriever. Meanwhile, the downstream selector can only improve knowledge precision, but it is likely to also reduce recall. Therefore, we can visualize the effect of applying a knowledge selector: it moves the RAG performance down and to the right in Figure 2; the better the knowledge selector, the more it will move right (improving precision), and the less it will move down (reducing recall).
@@
We find that for strong generator models, the knowledge recall score is the best single knowledge metric for estimating generation performance; Figure 3 shows a very strong correlation between knowledge recall and answer F1 for GPT-4o-mini on HotpotQA, and Figure 11 in the Appendix shows the same for LLaMA 3.1 8B on HotpotQA and for both models on WoW. Moreover, knowledge recall is the most important factor in improving generation performance given a fixed generator, while the knowledge precision score is a secondary factor.
@@
Figure 4: Color contours of answer F1 versus knowledge precision for GPT-4o-mini (top), LLaMA 3.1 8B (middle), and Mistral 7B-Instruct (bottom); the left column shows results on WoW, and the right shows HotpotQA. Each contour represents a different knowledge recall score; moving left to right visualizes improving the performance (precision) of the knowledge selector. Each plot point corresponds to a full experiment on the entire sampled dataset.
@@
All of our results show that the overall RAG performance depends on the generator model: the cross-model comparisons in Figures 2 & 4 show that for stronger generator models (as measured by rank in LLM leaderboards***https://huggingface.co/spaces/lmsys/chatbot-arena-leaderboard), performance across all knowledge settings are higher (Tables 1 & 2). While this result is unsurprising, we also find that the gap between the “full knowledge” and “gold knowledge” settings becomes narrower, suggesting that stronger generators are more robust to noisy input knowledge and rely less on the knowledge selector. In contrast, when the generator is weaker, any reasonable knowledge selector becomes beneficial, likely because a weak generator cannot handle noisy input and requires a selector to filter out distractor knowledge — knowledge F1 best correlates with answer F1 for Mistral-7B-Instruct in Figure 3.
@@
In addition, Figure 4 shows that attempting to improve the knowledge selector on top of a weak retriever (i.e. increasing precision when recall is low) hurts generation performance on WoW, as well as for GPT-4o-mini on Hotpot QA. In this scenario, more total knowledge, regardless of noise, improves response generation. These counter-intuitive observations are likely due to the nature of the task and the annotation quality. The solution space in HotpotQA is small given a specific question, and HotpotQA has a cleaner separation between gold and distractor knowledge, whereas in WoW, there are more plausible responses given the same conversation history and knowledge, as well as “distractor" knowledge that may actually be relevant to the gold response (see Appendix A.4 for more analysis).
@@
Interestingly, we observe that the boundary of where the knowledge selector improves generation performance (the border between the orange and white areas in Figure 2) is convex for all three LLMs on WoW. This phenomenon can also be seen in Figure 4, where some color contours intersect with the “full knowledge answer F1” baseline (dashed red line) multiple times. In other words, we observe that generation performance is non-monotonic as knowledge precision increases with a fixed knowledge recall. The fact that this phenomenon is only observed on WoW may be due to the relatively noisy gold knowledge annotations in WoW; we can produce similar behavior by artificially injecting noise into HotpotQA’s gold knowledge annotations (Appendix A.5).
Constraining the knowledge size does not change generation accuracy.
One important motivation for using a knowledge selector is to reduce the total input length to the generator model. However, while computational costs can be reduced by only using the top-$k$ knowledge, we find that the overall relationship between the knowledge precision-recall and generation F1 observed in our simulations remains unchanged, regardless of the value of $k$ (Appendix A.6).